OpenStudy (anonymous):

What are the approximate solutions of 4x2 + 3 = -12x to the nearest hundredth? (1 point) x ≈ -3.23 and x ≈ 0.23 x ≈ -2.72 and x ≈ -0.28 x ≈ 0.28 and x ≈ 2.72 x ≈ -0.23 and x ≈ 3.23

OpenStudy (anonymous):

Rewrite the equation as \(4x^2 + 12x + 3 = 0\). If \(x\) is positive, then every term of the left-hand side is positive, so the left-hand side is positive and not equal to zero. This means that there are no positive solutions, so we can rule out any answers which suggest that \(x\) is positive. This gives us answer #2.

OpenStudy (anonymous):

Thank you so much! Im on a IEP program and im supposed to have help from my teacher, but she is never there to help!.

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

Mind answering another one?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve for x: 16x2 - 81 = 0 (2 points) x = 4, 9 x = -4, -9 x = 9/4 , x = - 9/4 x = 81/16, x = - 81/16

OpenStudy (anonymous):

Rearrange terms (using addition) so that the x^2 term is on its own on one side and the constant term on the other. Then divide by the coefficient of x^2. Finally, take square roots.

OpenStudy (anonymous):

I really dont know how to do that :/

OpenStudy (anonymous):

Maybe you should show me how far you got.

OpenStudy (anonymous):

x: 16x2 - 81 = 0 I think you divide 16 on both sides?

OpenStudy (anonymous):

Sure.